# Integrated modelling of equilibrium and transport in axisymmetric magnetic mirror fusion devices

### v1

@@ 4.2 Parametric simulations of end plug performance
$B_{0}$ 4.0 T 5.0 T
$\langle n_{p0}\rangle$ $1.66\times 10^{20}$ m-3 $1.35\times 10^{20}$ m-3
$\langle\beta_{p0}\rangle$ 0.58 0.53
$E_{\mathrm{NBI}}$ 240 keV 360 keV
@@
$B_{0c}$ 3.0 T 2.6 T
$\beta_{c}$ 0.6 0.6
$n_{c}$ $7.5\times 10^{19}$ m-3 $5.8\times 10^{19}$ m-3
@@
Table 3: Optimized tandem parameters based on the results of parametric scans and POPCON analysis using the end plug conditions obtained in the scans.
While the optimized plugs identified in this scan are sufficient to deliver a viable central cell, this scan was coarse. It is foreseeable that performance, i.e. $\langle n\rangle/P_{\mathrm{NBI}}$, could be improved further by using Bayesian optimization coupled with a technique like Gaussian process machine learning. Use of such an approach would facilitate more efficient sampling of points in the multidimensional optimization space ($E_{\mathrm{NBI}}$, $P_{\mathrm{NBI}}$, $a_{m}$, $B_{m}$, $B_{0}$, $\ell$) such that a more optimized operating point could be found. It also could allow expansion of the optimization space to include parameters such as detailed aspects of the NBI injection geometry (focal length, source size, etc.), and optimization of the NBI beamline interference versus the real magnet geometry. The scalable nature of the IPS driven simulations here should facilitate such optimization procedures, and similar optimization procedures have been successful when applied to various aspects of tokamak design in the past (Nunn et al., 2023; Pavone et al., 2023; Brown et al., 2024; Rodriguez-Fernandez et al., 2024).
@@ 5 Conclusions and future work
Here, a new computational model for magnetic equilibrium, transport, heating, and fueling in simple mirrors was developed, a novel POPCON technique for tandem mirror performance was derived and optimized end plugs that deliver tandem mirror central cell performance consistent with the fusion pilot plant criteria outlined in the NASEM recommendations in the "Bringing Fusion to the US Grid" report were identified. The analysis here has identified an optimized $25~{}\mathrm{T}$ peak field tandem mirror end plug that achieves average densities in excess of $1.5\times 10^{20}~{}\mathrm{m}^{-3}$ with $15~{}\mathrm{MW}$ of $240-360~{}\mathrm{keV}$ NBI heating and enables a $50~{}\mathrm{m}$ long $3~{}\mathrm{T}$ central cell to generate over $150~{}\mathrm{MW}$ fusion power. This operating point is significantly more conservative than previous classical tandem mirror concepts proposed in the literature (Fowler & Logan, 1977; Fowler et al., 2017). Along the way, a number of important new and important insights about classical axisymmetric tandem mirror operation were brought to light. Using the novel POPCONs approach derived here, it was found found that in a classical tandem mirror the end plug can be characterized sufficiently to estimate central cell performance with only parameters: $a_{m}$ the plasma radius at the mirror throat, $B_{m}$ the field at the mirror throat, and $n_{p}$ the end plug density. The POPCONs found that alpha particle heating in the central cell can substantially raise $T_{e}$ and is important to understanding tandem mirror performance, and the central cell density ratio $n_{c}/n_{p}$ in high-field axisymmetric tandem mirrors should be significantly higher than previously suggested in the literature. End plug performance and engineering feasibility were also found to be enhanced by fueling the end plug with only tritium. Finally, first of a kind optimizations with integrated modelling for simple mirrors was performed using the RealTwin and it was shown that pilot plant end plug parameters can be reached.
@@
Along with these new discoveries, this work has exposed a number of important computational and theoretical capability gaps which need to be filled to further the modelling of tandem mirrors. Most importantly, the models here used a piece-wise approach to estimate central cell performance based on end plug performance. This required a number of assumptions. Ideally, a fully integrated model of the tandem mirror would be developed in the future. This could entail either a 2 real-space dimension 2 velocity-space dimension (2D2V) time-dependent Fokker-Planck solver or a set of bounce-averaged 1D2V time dependent Fokker-Planck solvers coupled together with the proper boundary-conditions and source/sink terms. A 2D2V FP solver would have the added benefit of resolving many of the issues which can arise with the ambipolar potential calculation on the addition of EC heating as a more accurate method for calculating the ambipolar potential such as that proposed in Degond et al. (2010) or Kupfer et al. (1996) could be used. Such a solver would also allow the investigation of thermal barrier tandems, which could provide a large performance enhancement over the classical tandems investigated here. While it was found that the applicability of an RF ion acceleration scheme analogous to that in WHAM is limited in an end plug (see Appendix B), further work on RF heating in magnetic mirrors is warranted. This would include the development of axisymmetric mirror geometry full-wave solvers and more innovative RF schemes which reduce quasilinear particle losses. Finally, while of limited importance in a negative-ion neutral beam fueled end plug, the impact of halo neutrals generated by NBI charge exchange and edge neutral charge exchange induced loss of fast ions are anticipated to be of some importance in smaller mirrors like WHAM. To better understand these effects the modelling tools used in this study will be integrated with edge neutral modelling codes.

### v2

@@ 4.2 Parametric simulations of standalone simple mirror performance
$B_{0}$ 4.0 - 6.0 T 5.5 - 9.0 T
$E_{\mathrm{NBI}}$ 240 - 480 keV 250 - 500 keV
$P_{\mathrm{NBI}}$ 10 - 20 MW 10 - 20 MW
Table 1: Mirror parameter ranges used in the parametric scan in Section04.2 and in the machine learning aided optimization in Section04.3.
With the constraints in the previous section established, a parametric scan in $B_{0}$, $P_{\mathrm{NBI}}$, $E_{\mathrm{NBI}}$ and $a_{m}$ was carried out with the RealTwin model described in Section 2 was specified. This scan used the end plug parameters shown in Table 1 to simulate a standalone simple mirror. For this work, points in the scan were selected over a linear range for each of the parameters. This scoping method is coarse, but substantial constraints were already placed on the problem through the use of the POPCON analysis in Section 3 and the analytic arguments used to constrain the physics and engineering earlier in this section. This made the coarse high fidelity scan sufficient to deliver an optimized simple mirror that exceeded the end plug requirements identified by the POPCON analysis. A scan in $\ell$ was also performed to determine if scaling power density with $\ell$ versus $P_{\mathrm{NBI}}$ was substantially different (changing the aspect ratio of the plug can slightly change NBI path length and modify NBI absorption). Varying $\ell$, however, had effectively the same impact on power density and plug performance as varying $P_{\mathrm{NBI}}$. In the rest of the discussion here will focus on the highest performing $\ell=4.5~{}\mathrm{m}$ case.
The CQL3D-m simulations performed in this section used 24 flux surfaces evenly spaced from 0.01 to 0.9 in normalized square-root poloidal flux $\sqrt{\psi_{n}}$, and the axial $z$ grid used to formulate the ambipolar response along the field line had 64 evenly spaced points extending from the midplane to the mirror throat. The distribution functions were discretized over 500 points in momentum-per-rest-mass $u_{0}$ with a grid which used increased resolution at low energies as to accurate capture the ion distributions and 400 evenly spaced points in the pitch angle $\vartheta$. The maximum energy in the momentum-per-rest-mass grid was set to correspond to a $250~{}\mathrm{keV}$ electron energy. The CQL3D-m simulations used 800 timesteps. The time advance was initially set to $1~{}\mathrm{ms}$ for the first 200 time steps, then increased to $2~{}\mathrm{ms}$ for the next 200 time steps, and then further increased to $4~{}\mathrm{ms}$ for the remaining 400 timesteps providing an overall simulation runtime of $2.2~{}\mathrm{s}$, several $\tau_{p}$. Ramped time-stepping was used to maintain numerical stability, as despite being an implicit time-advance code, the iterations of CQL3D-m with Freya and Pleiades and the internal ambipolar potential solve are explicit. The Freya NBI deposition calculation was run at each CQL3D-m time step using $2.5\times 10^{6}$ particles, and the Pleiades equilibrium was updated after every 100 time steps.
The results of the simulations using the RealTwin model are shown in Figure 8. Simulations which failed to satisfy the DCLC constraint (35b) or in which the equilibrium failed to converge due to $\beta>1$ or the onset of the mirror instability were removed from this analysis (this is why there is an empty regions in the Figure 8 plots. The onset of the mirror instability as the source of a $\beta$ limit in mirror plasmsas will be discussed further in a future publication). A potentially surprising result of these scans is that there is little difference in peak density versus $E_{\mathrm{NBI}}$, however, remembering the analysis in Section 4.1, this makes sense. Neutral beam absorption drops with higher $E_{\mathrm{NBI}}$, but the NBI fueling efficiency increases $\propto E_{\mathrm{NBI}}^{1/2}$. Another observation is that $\langle n\rangle$ improves superlinearly with increasing $P_{\mathrm{NBI}}$. This is due to a twofold non-linearity in the NBI absorption equation (35a) with NBI power. As $P_{\mathrm{NBI}}$ increases, the $\langle n\rangle$ increases and therefore the neutral beam absorption fraction increases. In addition, the plasma $\beta$ also increases as $P_{\mathrm{NBI}}$ increases. This causes the plasma to expand outwards from diagmagnetism, increasing the NBI absorption path length, further increasing NBI absorption.
Figure 9: Density $n_{p}$ and average ion energy $\langle E_{i}\rangle$, electron energy $\langle E_{e}\rangle\equiv 1.5T_{e}$, and ambipolar potential $e\phi$, versus the square root of the normalized poloidal flux $\sqrt{\psi_{n}}$. The profiles come from a simulation with parameters: $a_{m}=0.15~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=4~{}\mathrm{T}$, $\ell=4.5~{}\mathrm{m}$, $E_{\mathrm{NBI}}=240~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$.
As anticipated in Section 4.1, the target end plug $\langle n\rangle$ values identified in Section 3 were achieved at the lowest $P_{\mathrm{NBI}}$ for $E_{\mathrm{NBI}}$ of $240~{}\mathrm{keV}$ and $360~{}\mathrm{keV}$. Sufficient end plug conditions for fusion relevant central cell operation were achievable with $B_{0}$ of $4-6~{}\mathrm{T}$, $P_{\mathrm{NBI}}\sim 15~{}\mathrm{MW}$, $E_{\mathrm{NBI}}\sim 240-360~{}\mathrm{keV}$, and $a_{m}$ of $0.15-0.2~{}\mathrm{m}$. Under these conditions effectively full NBI absorption (>95%) was obtained with a $240~{}\mathrm{keV}$ beam, so lowering NBI energy further would only serve to reduce performance. The plasma profiles of the most optimal mirror identified in this scan are shown in Figure 9. The mirror had parameters: $a_{m}=0.15~{}\mathrm{m}$, $B_{0}=4~{}\mathrm{T}$, $E_{\mathrm{NBI}}=240~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}=15~{}\mathrm{MW}$ and achieved $\langle n\rangle>1.5\times 10^{20}~{}\mathrm{m}^{-3}$, more than sufficient to drive a $50~{}\mathrm{m}$ long central cell to $Q>5$ conditions based on the analysis in Section 3.
Parameter Optimum Standalone Optimum Consistent $T_{e}$
End plug
Ion Species T D
$a_{m}$ 0.15 m 0.25 m
$a_{0}$ 0.47 m 0.59 m
$\ell_{p}$ 4.5 m 4.5 m
$B_{m}$ 25 T 25 T
$B_{0p}$ 4.0 T 6.0 T
$\langle n_{p0}\rangle$ $1.66\times 10^{20}$ m-3 $1.60\times 10^{20}$ m-3
$\langle\beta_{p0}\rangle$ 0.58 0.59
$E_{\mathrm{NBI}}$ 240 keV 403 keV
$T_{e}$ 11 keV 120 keV
@@
$\theta_{\mathrm{NBI}}$ $45^{\circ}$ $50^{\circ}$
$\langle a_{0}/\rho_{0i}\rangle$ 24 29
@@
$B_{0c}$ 3.125 T 3.125 T
$\beta_{c}$ 0.6 0.55
$n_{c}$ $7.5\times 10^{19}$ m-3 $7.0\times 10^{19}$ m-3
@@
Table 2: Optimized tandem mirror parameters based on the results of parametric scans and POPCON analysis using the end plug conditions obtained using the standalone simulations with CQL3D-m and Pleiades performed in Section04.2 in column "Optimum Standalone" and the CQL3D-m and Pleiades simulations with electron temperature fixed to the central cell value performed in Section04.3 in column "Optimum Consistent $T_{e}$".
4.3 Optimization for a self-consistent end plug operating condition
While the end plug-like simple mirror identified in Section 4.2 appears to be sufficient to deliver a viable central cell, the scan those conditions were derived from was coarse. It also did not include the effects of increased electron temperature in the end plug as a result of alpha particle heating from the central cell. To address this, a secondary set of RealTwin simulations were created. These simulations used a modified CQL3D-m model in an attempt to better capture the tandem mirror physics. The electron distribution was assumed to be fixed by the classical tandem mirror central cell dynamics and the CQL3D-m simulations in this section used a fixed $T_{e}$ Maxwellian electron background where $n_{e}=n_{i}$ and a fixed ambipolar potential $\phi_{e}$ corresponding to those computed in the tandem mirror POPCON using (15) and (22). The operating point for the tandem mirror was chosen to be at a thermally stable ignited central cell (i.e. operating on the right hand side of the ignited region) operating point with $\beta_{c}\sim 0.6$ with the same plug density as the RealTwin simulations. The POPCON and the RealTwin simulations were iterated until a self-consistent condition where $T_{e}$, $\phi_{e}$, and $\langle n_{p}\rangle$ in the POPCON and RealTwin models matched. These RealTwin simulations consistent with the POPCON conditions provide a more detailed picture of steady-state end plug operation in a tandem mirror. Most notably, the larger ambipolar hole in the ion distribution increases the optimal values of $E_{\mathrm{NBI}}$ and $B_{0}$.
In addition to fixed $T_{e}$ and $\phi_{e}$, these simulations employed several other modifications and refinements on the model versus Section 4.2. Rather than the coarse parameteric scan used in the previous section, the Bayesian optimization technique described in Section 2 was used to efficiently find values of $a_{m},$ $B_{0},$ $E_{\mathrm{NBI}},$ and $P_{\mathrm{NBI}}$ which optimize the normalized Lawson criterion $\langle n\tau\rangle(P_{\mathrm{abs}}/P_{\mathrm{NBI}})$, where $P_{\mathrm{abs}}/P_{\mathrm{NBI}}$ is the absorbed fraction of the total NBI power. This metric can be thought of as a measure of NBI fuelling efficiency. The normalized Lawson criteria is the preferred objective function quantity as it simultaneously optimizes density per unit of absorbed NBI power and ensures good beam absorption. Rather than the simple $a_{0}/\rho_{i0}>25$ DCLC constraint used previously, the $\beta$ adjusted DCLC constraint from Tang et al. (1972)
$N_{\rho}^{\mathrm{Tang}}\equiv a_{0}/\rho_{i0}\approx\frac{1}{2}\frac{1}{\sqrt{\beta}\sqrt{m_{e}/m_{i}+\Omega^{2}_{i}/\omega_{pi}^{2}}},$ (38)
was employed in order to provide a more accurate estimate of the DCLC stability. This constraint was implemented automatically in the objective function through the use of penalty function: $f(a_{0}/\rho_{i})=1/2\tanh{\{[a_{0}/\rho_{i}-(N_{\rho}^{\mathrm{Tang}}-5)]/5\}}$ which multiplies $\langle n\tau\rangle(P_{\mathrm{abs}}/P_{\mathrm{NBI}})$.
The optimization in this section was initially performed using tritium to improve classical confinement. However, the higher $E_{\mathrm{NBI}}$ required to overcome the ambipolar hole at high $T_{e}$ caused end plugs using tritium as the main ion species to not satisfy the DCLC constraint. Because of this, a switch to deuterium was made to reduce $\rho_{i}$. The confinement reduction factor resultant from this switch was $\sim\sqrt{3/2}$ rather than $\sim 3/2$ as the elevated $T_{e}$ in these simulations caused the plasma to be ion-ion scattering dominated rather than electron drag dominated as it was in Section 4.2.
In order to recover some classical confinement performance in deuterium operation, the NBI angle $\theta_{\mathrm{NBI}}$ was shifted from the value of $45^{\circ}$ used in Section 4.2 to $50^{\circ}$, increasing the effective mirror ratio. This small shift was acceptable as $\theta_{\mathrm{NBI}}$ should not be perpendicular enough to excite the Alfven ion-cyclotron instability (Smith, 1984) or the mirror instability (14), sloshing ions are not required for DCLC stability, and the NBI path length reduction is small enough that the shift in $\theta_{\mathrm{NBI}}$ does not significantly reduce ${P_{abs}}/{P_{\mathrm{NBI}}}$.
Figure 10: Plots of density $n$ (blue) and ion temperature $E_{i}$ (red) versus normalized square root poloidal flux $\sqrt{\psi_{n}}$ obtained from CQL3D-m simulations of a fixed $\beta=0.6$ end plug with $a_{m}=0.25~{}\mathrm{m}$, $B_{m}=25~{}\mathrm{T}$, $B_{0}=6~{}\mathrm{T}$, $E_{\mathrm{NBI}}=400~{}\mathrm{keV}$, $P_{\mathrm{NBI}}=20~{}\mathrm{MW}$, and $T_{e}=120~{}\mathrm{keV}$ both with (solid) and without (dashed) classical radial diffusion. Radial diffusion has a significant affect on confinement in these cases. Diffusion notably reduces the total density in addition to flattening the density. The $E_{i}>E_{\mathrm{NBI}}$ seen in the plot is the result of a combination of up-scattering in ion energy and the preferential loss of low energy ions.
End plug simulations, especially those with large $E_{\mathrm{NBI}}$, were found to exhibit an instability tied to $\beta\rightarrow 1$ near the magnetic axis. The higher NBI power density due to smaller flux surface volumes as $\sqrt{\psi_{n}}\rightarrow 0$ (low $\sqrt{\psi_{n}}$ surfaces absorb a similar amount of NBI power to large $\sqrt{\psi_{n}}$ surfaces for an NBI width substantially less than $a_{0}$) caused $\beta$ on the inner flux surfaces to grow quickly. Higher $\beta$ increased $R_{m}$ and $\tau_{p}$. This in turn increased $n$ leading to a further increase to the NBI power density on the inner flux surfaces. This feedback loop rapidly caused $\beta$ on the inner flux surfaces to exceed unity resulting in a loss of magnetic equilibrium. To combat the instability, a radially varying classical diffusion coefficient based on (21) was added to CQL3D-m’s Fokker-Planck equation in order to flatten the pressure profiles in the end plug. A typical value for this diffusion coefficient is $D_{\rho}\sim 1\times 10^{-4}~{}\mathrm{m^{2}s^{-1}}$. This addition caused a reduction in performance due to the introduction of radial losses, but classical diffusion successfully resolved the $\beta\rightarrow 1$ instability in most cases and flattened the radial end plug density profiles. An example of the profile flattening from radial diffusion is shown in Figure 10.
However, when the magnetic equilibrium was evolved self-consistently with Pleiades using the pressure profiles from the CQL3D-m simulations with diffusion, numerical noise developed which could corrupt the solution. The noise appeared to be related to the radial boundary conditions used in CQL3D-m, and there will be attempts to resolve this noise in future work so that fully self-consistent CQL3D-m/Pleiades optimizations of end plug performance may be performed. To combat the noise here, fixed $\beta_{p}=0.6$ Pleiades equilibria were used during the optimization to prevent failures which can impair the GPBO’s convergence. To maintain a degree of self-consistency, the optimization imposed a Gaussian envelope penalty function (centred at $\beta_{p}=0.6$ with $\sigma=0.1$) multiplying $\langle n\tau\rangle(P_{\mathrm{abs}}/P_{\mathrm{NBI}})$ to ensure that optimized end plugs had $\beta_{p}\sim 0.6$. After optimal end plug conditions were identified, the optimal cases were rerun in standalone simulations which self-consistently included field evolution to confirm that the same results could be obtained.
The CQL3D-m simulations in this section could be evolved for a longer period of time at lower resolution as the explicit potential solver was disabled and electrons were considered as a background species rather than undergoing nonlinear evolution with the ions. These CQL3D-m simulations used 24 flux surfaces in the radial grid even spaced from $\sqrt{\psi_{n}}=0.01-0.95$ and 64 axial points evenly spaced in $\hat{z}$ between the midplane and the mirror throat. The distribution function was discretized over 300 points in momentum-per-rest-mass $u_{0}$ and 200 evenly spaced points in the pitch angle $\vartheta$. The maximum energy in the momentum-per-rest-mass grid was set to correspond to $1000~{}\mathrm{keV}$ deuterons. CQL3D-m used 100 ms timesteps and Freya NBI deposition calculations were run at each time step using $2.5\times 10^{6}$ particles. Simulations were evolved for $10~{}\mathrm{s}$, several $\tau_{p}$, to time converge the end plug transport. Rather than self-consistently optimize both the POPCON and the end plug in a single simulation loop, optimal end plug conditions were calculated at a number of different values of $T_{e}$ independently using the Bayesian optimization procedure described in Section 2. At each value of $T_{e}$, five batches of 10 simulations were run with parameters selected using the CL-min acquisition function. The initial three batches used a more exploratory acquisition function with $\kappa=10$ and final two batches used a lower $\kappa=5$ which allowed the optimization to quickly find a global optimum.
Figure 11: A POPCON of the tandem mirror operating point described the second column labelled "Optimum Consistent $T_{e}$" Table 2. Heating power $P_{\mathrm{RF}}$ is shown in the filled red (positive) and blue (negative, ignited) contours, fusion power $P_{fus}$ in black contours, $\beta_{c}$ in green contours, and the operating point in the table is marked with a star.
As anticipated in Section 4.2, when self-consistent electron temperatures were accounted for, $E_{\mathrm{NBI}}$, $B_{0}$, and $a_{m}$ had to increase to ensure that the plasma remained below the $\beta$ limit, had good confinement, had adequate NBI absorption, and remained stable to DCLC. The optimization procedure selected the largest end plug radii $a_{m}=25~{}\textrm{cm}$ set by engineering feasibility as it both improved $\langle n\tau_{p}\rangle(P_{\mathrm{abs}}/P_{\mathrm{NBI}})$ and allowed the plasma to more easily satisfy the DCLC stability condition (38). With end plug $T_{e}=80-120~{}\mathrm{keV}$, $E_{\mathrm{NBI}}\approx 400~{}\mathrm{keV}$, and $P_{\mathrm{NBI}}\approx 15-20~{}\mathrm{MW}$, desirable $1.5\times 10^{20}~{}\mathrm{m}^{-3}$ densities were achievable. These densities were achievable despite the $a_{m}=25~{}\mathrm{cm}$ plugs’ higher volume because of the increased fuelling efficiency obtained with high-energy NBI and the reduction in electron drag from using self-consistent values for $T_{e}$. The best end plug performance was obtained at $T_{e}=100~{}\mathrm{keV}$ as it provided a good balance of increased ambipolar losses and decreased electron drag. However, for the best overall system performance occurred at $T_{e}=120~{}\mathrm{keV}$, and it was chosen for the optimized operating point as it increased the achievable central cell $P_{fus}$ by $\sim 100~{}\mathrm{MW}$. The optimized tandem mirror POPCON operating point based on this optimization process is shown in Figure 11 and the parameters of the optimized tandem mirror are given in the second column labelled "Optimum Consistent $T_{e}$" Table 2. It is noted that, in future evolutions of this design, slightly higher $\beta$ may be desirable in order to achieve high $\beta$ stabilization using conducting walls (Berk et al., 1987; Li et al., 1987) and higher $P_{\mathrm{fus}}$. This may be achieved under the same parameters with elevated $P_{\mathrm{NBI}}$, or can be optimized for in a new optimization. The optimized tandem mirror operating point in Table 2 has self-consistent $T_{e}$, uses technologically feasible components, obeys the DCLC stability constraint (38) (the Tang constraint specified $a_{0}/\rho_{i}>27$ for diamagnetic drift wave stability), and satisfies the NASEM requirements for a fusion power core (Hawryluk et al., 2021).
@@ 5 Conclusions and future work
A new computational model for magnetic equilibrium, transport, heating, and fuelling in simple mirrors was developed, a novel POPCON technique for tandem mirror performance was derived, and optimized end plugs that deliver tandem mirror central cell performance consistent with the fusion pilot plant criteria outlined in the NASEM recommendations in the “Bringing Fusion to the US Grid” report were identified. The analysis here has identified an optimized $25~{}\mathrm{T}$ peak field tandem mirror end plug that achieves average densities in excess of $1.5\times 10^{20}~{}\mathrm{m}^{-3}$ with $15~{}\mathrm{MW}$ of $240-360~{}\mathrm{keV}$ NBI heating and enables a $50~{}\mathrm{m}$ long $3~{}\mathrm{T}$ central cell to generate over $150~{}\mathrm{MW}$ fusion power. This operating point is significantly more conservative than previous classical tandem mirror concepts proposed in the literature (Fowler & Logan, 1977; Fowler et al., 2017). Along the way, a number of important new and important insights about classical axisymmetric tandem mirror operation were brought to light. Using the novel POPCONs approach derived here, it was found found that in a classical tandem mirror the end plug can be characterized sufficiently to estimate central cell performance with only parameters: the plasma radius at the mirror throat $a_{m}$, the field at the mirror throat $B_{m}$, and the end plug density $n_{p}$. The POPCONs found that alpha particle heating in the central cell can substantially raise $T_{e}$ and is important to understanding tandem mirror performance, and the central cell density ratio $n_{c}/n_{p}$ in high-field axisymmetric tandem mirrors should be significantly higher than previously suggested in the literature. It was also noted that synchrotron radiation and classical transport could substantially limit tandem mirror central cell performance when $\beta$ becomes large, and therefore these effects must be included in the tandem mirror scopings. End plug performance and engineering feasibility were also found to be enhanced by fuelling the end plug with only tritium provided the DCLC stability conditions could be satisfied. Machine learning aided optimization of end plug equilibria and transport using integrated modelling was performed using Realta’s RealTwin integrated model. It was shown that pilot plant relevant end plug parameters can be reached with near-term technologies. These simulations also revealed two important pieces of end plug physics that have been overlooked in previous studies such as Fowler et al. (2017). First, NBI energy optimizes at a significantly lower value, $\sim 400~{}\textrm{keV}$, than what was used in previous tandem mirror scoping studies, $\sim 1~{}\textrm{MeV}$, when neutral beam shine through and DCLC are self-consistently accounted for. Second, classical transport is very important to moderating the value of $\beta_{0}$ at the magnetic axis of the end plug. Without diffusion, $\beta_{0}(\psi=0)$ in the end plug will rapidly approach unity for realistic NBI absorption profiles limiting the maximum value of $\langle n_{p}\rangle$. Furthermore, the flattening of end plug density profiles from diffusion could suppress drift waves, having a stabilizing effect on DCLC.
@@
Along with these new discoveries, this work has exposed a number of important computational and theoretical capability gaps which need to be filled to further the modelling of tandem mirrors the most obvious of which is kinetic and MHD stability which were not considered rigorously here. Most stability actuators, such as shear flow, entail nonlinear stabilization processes that are difficult to resolve in stimulation and will require new nonlinear stability models like that shown in Tran et al. (2024) coupled with the equilibrium models presented here. In addition to the limitations associated with predicting stability the equilibrium models here the piece-wise approach to estimate central cell performance based on end plug performance required a number of assumptions. Ideally, a fully integrated model of the tandem mirror transport would be developed in the future. This could entail either a 2 real-space dimension 2 velocity-space dimension (2D2V) time-dependent Fokker-Planck solver or a set of bounce-averaged 1D2V time dependent Fokker-Planck solvers coupled together with the proper boundary-conditions and source/sink terms. A 2D2V FP solver would have the added benefit of resolving many of the issues which can arise with the ambipolar potential calculation on the addition of EC heating as a more accurate method for calculating the ambipolar potential such as that proposed in Degond et al. (2010) or Kupfer et al. (1996) could be used. Such a solver would also allow accurate consideration magnetic expander regions and the investigation of thermal barrier tandems, which could provide a large performance enhancement over the classical tandems investigated here. While it was found that the applicability of an RF ion acceleration scheme analogous to that in WHAM is limited in an end plug (see Appendix A), further work on RF heating in magnetic mirrors is warranted. This would include the development of axisymmetric mirror geometry full-wave solvers and more innovative RF schemes which reduce quasilinear particle losses. Finally, while of limited importance in a negative-ion neutral beam fueled end plug, the impact of halo neutrals generated by NBI charge exchange and edge neutral charge exchange induced loss of fast ions are anticipated to be of some importance in smaller mirrors like WHAM. To better understand these effects the modelling tools used in this study will be integrated with edge neutral modelling codes.